# Depth-Breadth Synergy in RLVR: Unlocking LLM Reasoning Gains with Adaptive Exploration

### v1

@@ 5.2 Analysis of RLVR
With the rapid advancement of RLVR and the proliferation of open-source LRMs, many studies have begun to analyze the RLVR pipeline and these open LRMs. Several studies (Liu et al., 2025a; Zhao et al., 2025; Shah et al., 2025) indicates that the self-reflect and self-critique behaviors observed after RLVR originates from the base model rather than the RL process. Dang et al. (2025) find that although the RLVR process benefits Pass@1, Pass@K may decline as training progresses. Subsequently, Yue et al. (2025a) through extensive experimental analysis, discovered that RLVR’s performance is significantly constrained by the base model; once training converges, it struggles to surpass the capability boundary of the base model. These studies have sparked widespread concern about the capability ceiling of RLVR, and consequently, the Pass@K metric has become a focal point for diagnosing and potentially transcending the intrinsic limits imposed by the base model. This paper analyzes and refines the RLVR pipeline from the dual perspectives of Pass@1 and Pass@K.
@@ 6 Conclusion
In this work, we identify and systematically analyze two critical yet under-explored dimensions in Reinforcement Learning with Verifiable Reward (RLVR): Depth (the hardest problems a model can learn) and Breadth (the number of instances processed per iteration). We reveal that existing GRPO-based methods suffer from a fundamental limitation—their cumulative advantage calculation disproportionately weights medium-difficulty problems while neglecting hard ones, creating a bottleneck for Pass@K performance. To address this issue, we propose Difficulty-Adaptive Rollout Sampling (DARS) to reallocate compute to the hard problems, boosting Pass@K with low extra inference cost. Furthermore, we demonstrate that large-breadth training acts as an orthogonal dimension for RLVR optimization. Scaling batch size and employing full-batch updates significantly boost Pass@1 performance while sustaining high token-level entropy. By unifying these dimensions in our DARS-B framework, we achieve simultaneous gains in both Pass@1 and Pass@K, confirming the synergistic relationship between depth and breadth in RLVR.
Appendix A Appendix

### v2

@@ 5.2 Analysis of RLVR
With the rapid advancement of RLVR and the proliferation of open-source LRMs, many studies have begun to analyze the RLVR pipeline and these open LRMs. Several studies (Liu et al., 2025a; Zhao et al., 2025; Shah et al., 2025) indicates that the self-reflect and self-critique behaviors observed after RLVR originates from the base model rather than the RL process. Dang et al. (2025) find that although the RLVR process benefits Pass@1, Pass@K may decline as training progresses. Subsequently, Yue et al. (2025a) through extensive experimental analysis, discovered that RLVR’s performance is significantly constrained by the base model; once training converges, it struggles to surpass the capability boundary of the base model. These studies have sparked widespread concern about the capability ceiling of RLVR, and consequently, the Pass@K metric has become a focal point for diagnosing and potentially transcending the intrinsic limits imposed by the base model (Liang et al., 2025). This paper analyzes and refines the RLVR pipeline from the dual perspectives of Pass@1 and Pass@K.
@@ 6 Conclusion
In this work, we identify and systematically analyze two critical yet under-explored dimensions in Reinforcement Learning with Verifiable Reward (RLVR): Depth (the hardest problems a model can learn) and Breadth (the number of instances processed per iteration). We reveal that existing GRPO-based methods suffer from a fundamental limitation—their cumulative advantage calculation disproportionately weights medium-difficulty problems while neglecting hard ones, creating a bottleneck for Pass@K performance. To address this issue, we propose Difficulty-Adaptive Rollout Sampling (DARS) to reallocate compute to the hard problems, boosting Pass@K with low extra inference cost. Furthermore, we demonstrate that large-breadth training acts as an orthogonal dimension for RLVR optimization. Scaling batch size and employing full-batch updates significantly boost Pass@1 performance while sustaining high token-level entropy. By unifying these dimensions in our DARS-Breadth framework, we achieve simultaneous gains in both Pass@1 and Pass@K, confirming the synergistic relationship between depth and breadth in RLVR.
Appendix
Appendix A Training and Evaluation Details
@@ A.1 Setting Details
Parameters and Metrics.
Currently, our experiments are conducted with Qwen2.5-Math series language models (Yang et al., 2024). We set the temperature to 1.0 for both the training and evaluation procedures. In this paper, we mainly use two metrics, Pass@1 and Pass@K. To acquire Pass@K results, we sample 128 candidate responses for each question during the evaluation process; the calculation of Pass@1 is derived from Avg@128. Both the training and evaluation processes are scored using Math-Verify. The learning rate is 1e-6 for depth training methods, and 5e-6 for large breadth training. We do not use the reference model and KL loss. For fair comparison, we uniformly set the PPO mini step to 2 for all experiments. By default, the maximum prompt length is 1024, and the maximum response length is 3072 for the Qwen2.5-Math series model.
Implementation Details.
Following LUFFY (Yan et al., 2025), we use the default subset and filter out generations that are longer than 8192 tokens and those that are verified wrong by Math-Verify  [^1] https://github.com/huggingface/Math-Verify, resulting in 45k question-solution pairs. For training Llama-3.1-8B, we use the train split of MATH dataset. Our training framework is derived from Verl (Sheng et al., 2024) pipeline, which is a flexible, high-performance reinforcement-learning framework built for training large language-model agents. With native PyTorch support and efficient distributed training, Verl lets researchers quickly prototype and scale RL algorithms like PPO on GPUs. All of our experiments are conducted on H200 GPUs. At present, the LLM of our experiment is the Qwen2.5-Math series.
Training Steps and Checkpoint Steps.
For non-breadth methods on Qwen2.5-Math-1.5B/7B, we set the checkpoint step as 100. For breadth methods on Qwen2.5-Math-1.5B/7B, we set the checkpoint step as 15. The specific training steps are determined according to the convergence of the model. The number of training steps for non-breadth training is set as 300 for Llama-3.1-8B, 600 for Qwen2.5-Math-1.5B, and 500 for Qwen2.5-Math-7B. The number of training steps for breadth training is set as 70. For breadth training, we set the total training steps as 105 for Qwen2.5-Math-1.5B, and 75 for Qwen2.5-Math-7B.
Appendix B More Experimental Results
B.1 Ablation Study on std-based Advantage Computation
As illustrated in Section 2, Dr. GRPO  (Liu et al., 2025b) removes the standard-deviation term from the advantage computation to eliminate question-level difficulty bias, and demonstrates its superiority through extensive experiments. Consequently, the experiments reported in this study were conducted primarily though the Dr. GRPO methodology. To further illustrate the effectiveness of DARS on std-based advantage computation, we conduct the experiment with HW schedule on Qwen2.5-Math-1.5B model, as shown in Figure 11
Figure 11: Comparison of our DARS on std-based advantage computation.
B.2 Depth and Breadth Synergy for Pass@1 and Pass@32
In Section 4.3, we show the training dynamics of Pass@128-Pass@1 for DARS and baseline methods. To further illustrate the effectiveness of DARS, we show the training dynamics of Pass@32-Pass@1 in Figure 12. Our DARS significantly improves the Pass@32 performance compared to other methods.
Figure 12: Complementary improve of Depth and Breadth Synergy for Pass@1 and Pass@K (K=32) performance.
B.3 Comparison of ET/HW Schedule in Breadth Scaling
In addition, compared with the ET schedule, DARS-HW-Breadth significantly improves the model’s Pass@128 performance as shown in Figure 13. We consider this performance gain is due to the HW schedule placing greater emphasis on difficult samples.
Figure 13: Comparison of ET and HW schedule in breadth training of Qwen2.5-Math series.
Appendix C Discussion and Future Work
In this section, we analyze how hyperparameters $N$ and $N^{\max}$ control the shape of the cumulative advantage curve, and how this shape may influence training behavior. We further discuss how dynamically adjusting these parameters could enable a smooth transition from Pass@K-oriented to Pass@1-oriented training.
C.1 Hyperparameter Control of Cumulative Advantage Shape
We show the Cumulative Advantage shape of ET/HW schedule with $N=8$ in Figure 14. By continuously reducing the size of $N_{\text{max}}$, the curve will contract accordingly. When $N_{\text{max}}=N$, it is equivalent to the vanilla method without performing DARS.
Figure 14: Control the shape of Cumulative Advantage by adjusting the $N_{max}$ hyperparameter of DARS.
C.2 Potential Pass@K to Pass@1 Training Transition
The dynamic control of $N^{\max}$ suggests an intriguing training strategy: starting with a large $N^{\max}$ value to maximize Pass@K performance through extensive exploration of hard problems, then gradually reducing $N^{\max}$ throughout training to transition toward Pass@1 optimization.
This approach mirrors curriculum learning principles, where the training difficulty is progressively adjusted. Initially, the model benefits from the expanded solution space and diverse reasoning patterns discovered through heavy sampling on hard problems (high $N^{\max}$). As training progresses and the model’s capability matures, reducing $N^{\max}$ focuses the training on refining the most promising solution strategies, ultimately improving single-shot performance.
Future work will explore optimal annealing schedules for $N^{\max}$ and investigate whether this transition strategy can simultaneously maximize both Pass@1 and Pass@K performance, potentially overcoming the current limitations of RLVR training.